# Transcorrelated Theory with Pseudopotentials

### v1

@@ 4.4 Dissociation energies
Figure 8 shows energy differences between experimental and theoretical evaluations of the N2 dissociation curve. Energy differences are presented as a function of bond length. The experimental curve is taken from Ref. 39. The theoretical curves are evaluated with FCIQMC and MRCI methods. We used eCEPPS and did the calculations in AVDZ, AVTZ, and AVQZ basis sets using FCIQMC. We show results for FCIQMC and xTC-FCIQMC(PP-$2$). MRCI-F12 results are only evaluated in the AVQZ basis set. The theoretical results are shifted to overlap with experiment at $r=4.2$, a length which corresponds to the system of two isolated nitrogen atoms.
@@
Figure 8: N2 FCIQMC (black crosses) and xTC-FCIQMC(PP-$2$) (red circles) dissociation curves as a function of bond length with eCEPPS, computed in AVDZ (top row), AVTZ (middle row) and AVQZ (bottom row) basis sets. MRCI-F12 energies are evaluated in AVQZ basis set in all of the images, and both eCEPP (dashed black) and all-electron MRCI-F12 (dotted black) curves are shown. MRCI-F12 results are evaluated in the AVQZ basis set in all of the images. Results are presented as differences to experimental dissociation curve from Ref. 39, taking the overlap between theory and experiment to be at $r=4.2$The grey shaded region denotes chemical accuracy with respect to experiment. The dash-dotted vertical line points to the equilibrium bond length of N2 of $r=1.098$.
@@
Because of the strong correlations involved in dissociation of the N2, the Hartree-Fock wave function is a poor reference for the Jastrow factor optimization. Hence we used a method developed recently to tackle this problem 40. First, we took the $100$ most populated determinants from a non-transcorrelated FCIQMC calculation, and used them for constructing a trial wave function for VMC, and optimized the Jastrow factor with the VMC method before using it for preparing the transcorrelated Hamiltonian. In the subsequent xTC phase we used the reduced density matrix of the multideterminant wave function used in the optimisation.
@@
Figure 8 shows that FCIQMC does not converge with respect to basis-set-error with any basis set used at small bond lengths. xTC-FCIQMC(PP-$2$) still overestimate the equilibrium energy at triple-zeta basis, but has an excellent match in quadruple zeta-basis, a feature already seen with xTC coupled cluster methods and eCEPPs in Fig. 6. The xTC-FCIQMC(PP-$2$) method in quadruple zeta-basis obtains chemical accuracy everywhere except at small bond lengths and at $r=2.098$ , where, however, the stochastic error of the result overlaps with chemical accuracy regime.
@@
Figure 9 shows the dissociation error curve of F2, showing the difference between MRCI-F12 and xTC-FCIQMC(PP-$2$) theoretical methods and experiment. xTC-FCIQMC(PP-$2$) results are evaluated with eCEPPS and computed in AVDZ, AVTZ, and AVQZ basis sets. The MRCI-F12 results are evaluated in the AVQZ basis set. The experimental curve is taken from Ref. 41. Because the experimental data extends only to $r\sim 2.8$, the theoretical results are shifted to overlap with experiment at the equilibrium bond length of $r=1.4118$.
@@
Figure 9: F2 xTC-FCIQMC(PP-$2$) and MRCI-F12 energies as a function of bond length with eCEPPS, computed in AVDZ (top row), AVTZ (middle row) and AVQZ (bottom row) basis sets. MRCI-F12 energies are the same in all plots, and are evaluated in AVQZ basis set. Results are presented as differences to experimental dissociation curve from Ref. 41. The grey shaded region denotes the chemical accuracy with respect to experiment. The dotted vertical line points to the equilibrium bond length of F2 of $1.4112$.

### v2

@@ 4.5 Dissociation energies
Figure 9 shows energy differences between experimental and theoretical evaluations of the N2 dissociation curve. Energy differences are presented as a function of bond length. The experimental curve is taken from Ref. 40. The theoretical curves are evaluated with FCIQMC and MRCI methods. We used eCEPPS and did the calculations in AVDZ, AVTZ, and AVQZ basis sets using FCIQMC. We show results for FCIQMC and xTC-FCIQMC(PP-$2$). MRCI-F12 results are only evaluated in the AVQZ basis set. The theoretical results are shifted to overlap with experiment at $r=4.2$, a length which corresponds to the system of two isolated nitrogen atoms.
@@
Figure 9: N2 FCIQMC (black crosses) and xTC-FCIQMC(PP-$2$) (red circles) dissociation curves as a function of bond length with eCEPPS, computed in AVDZ (top row), AVTZ (middle row) and AVQZ (bottom row) basis sets. MRCI-F12 energies are evaluated in AVQZ basis set in all of the images, and both eCEPP (dashed black) and all-electron MRCI-F12 (dotted black) curves are shown. MRCI-F12 results are evaluated in the AVQZ basis set in all of the images. Results are presented as differences to experimental dissociation curve from Ref. 40, taking the overlap between theory and experiment to be at $r=4.2$The grey shaded region denotes chemical accuracy with respect to experiment. The dash-dotted vertical line points to the equilibrium bond length of N2 of $r=1.098$.
@@
Because of the strong correlations and change of spin state involved in dissociation of the N2, the Hartree-Fock wave function is a poor reference for the Jastrow factor optimization. Hence we used a method developed recently to tackle this problem 41. First, we took the $100$ most populated determinants from a non-transcorrelated FCIQMC calculation, and used them for constructing a trial wave function for VMC, and optimized the Jastrow factor with the VMC method before using it for preparing the transcorrelated Hamiltonian. In the subsequent xTC phase we used the reduced density matrix of the multideterminant wave function used in the optimisation.
@@
Figure 9 shows that the regular (non-transcorrelated) FCIQMC does not achieve chemical accuracy with respect to basis-set-error at near-equilibrium bond lengths. xTC-FCIQMC(PP-$2$) underestimates the equilibrium energy at triple-zeta basis by about 4mH, but has an excellent match in quadruple zeta-basis, a feature already seen with xTC coupled cluster methods and eCEPPs in Fig. 7. The xTC-FCIQMC(PP-$2$) method in quadruple zeta-basis obtains chemical accuracy everywhere except at the most compressed bond length (below 1Å), and at $r=2.098$ , where, however, the stochastic error of the result overlaps with chemical accuracy regime.
We tested the size-consistency by calculating the difference between the FCIQMC energies of two isolated nitrogen atoms and the energy of the N2 molecule at the longest bond distance studied, $\Delta E=2E_{\text{N}}-E_{\text{N}_{2}}(r=4.2\AA)$. This test was done in AVDZ basis, and resulted in $\Delta E=-1.5(2)$mHa. This small size-inconsistency can be traced to the use of the combined Jastrow treatment (see discussion after Eq. (6)) together with the xTC approximation, since the latter approximates the effect of the three-body interactions introduced by the commutators of the kinetic energy operator and the Jastrow factor. In the combined Jastrow approach the one-body terms are folded into 2-body Jastrow terms, which in turn leads to additional 3-body interactions (approximated in the xTC treatment). This combination leads to size-inconsistency. This size-consistency error can be entirely eliminated by using the separated Jastrow treatment (See Supplementary material for the separate treatment of PP commutators). We verified explicitly with further calculations that the xTC approximation, in the separate Jastrow treatment, does not incur a size-consistency error. Because we checked that in equilibrium geometries the separate and combined treatments yielded the same results, the overall error in this study due to the aforementioned size-inconsistency is on the order of 1-2mHa. The best workflow for future studies will require further investigation of Jastrow cutoffs and the use of the separate Jastrow treatment.
@@
Figure 10 shows the dissociation error curve of F2, showing the difference between MRCI-F12 and xTC-FCIQMC(PP-$2$) theoretical methods and experiment. xTC-FCIQMC(PP-$2$) results are evaluated with eCEPPS and computed in AVDZ, AVTZ, and AVQZ basis sets. The MRCI-F12 results are evaluated in the AVQZ basis set. The experimental curve is taken from Ref. 42. Because the experimental data extends only to $r\sim 2.8$, the theoretical results are shifted to overlap with experiment at the equilibrium bond length of $r=1.4118$.
@@
Figure 10: F2 xTC-FCIQMC(PP-$2$) and MRCI-F12 energies as a function of bond length with eCEPPS, computed in AVDZ (top row), AVTZ (middle row) and AVQZ (bottom row) basis sets. MRCI-F12 energies are the same in all plots, and are evaluated in AVQZ basis set. Results are presented as differences to experimental dissociation curve from Ref. 42. The grey shaded region denotes the chemical accuracy with respect to experiment. The dotted vertical line points to the equilibrium bond length of F2 of $1.4112$.
@@ 5 Conclusions
Supporting Information
Supporting Information. • Optimized Jastrow factor parameters for atoms and molecules studied in this work. • Derivations of pseudopotential commutator formulas (separate Jastrow treatment). • FCIDUMP files containing transcorrelated second-quantized Hamiltonians for Be in AVDZ basis set, with and without the PP commutator evaluations.